13 - 51
15 - 49
22 - 42
33 - 31
36 - 28
Input / output .. I need the Rule here thanks I am stuck

Found 2 solutions by jerryguo41, stanbon:
Answer by jerryguo41(197)

13 - 51
(+2)-(-2)
15 - 49
(+7)-(-7)
22 - 42
Starting at 1 - 63 you can add to input the same amount as you subtract from output.

Answer by stanbon(75887)

13 - 51
15 - 49
22 - 42
33 - 31
36 - 28
Input / output .. I need the Rule here thanks
----------------
The relation is linear.
----
Use two of the pairs to find the slope:
m = (49-51)/(15-13) = -1
-----
Form: y = mx + b
Substitute for y,x and m, then solve for "b":
51 = -1(13)+b
b = 64
-------
Equation:
y = -x+64
================
Cheers,
stan H.